OpenStudy (anonymous):

Solve the equation. Round the answer to the nearest tenth. 2.8x = 5.1 A 1.7 B. 1.8 C. 1.9 D. 18.2

OpenStudy (ivettef365):

divide 5.1/2.8 what do you get ?

OpenStudy (anonymous):

Well, first we divide both sides by 2.8 \[ 2.8x = 5.1 \quad \implies \quad x = \frac{5.1}{2.8} \] This fraction is annoying, but we can multiply both the numerator and the divisor by 10 to get rid of the decimal point: \[ x = \frac{5.1}{2.8} = \frac{51}{28} \] And now to calculate this using long division. Notice we need to calculate up to the second digit after the decimal point, since we are asked to round into tenths and 0.09 rounds to 0.1 for example..  So \(x = 1.82...\) and closest to that is B) 1.8 I know it's simple stuff, but if you happened to forgot the long division thing then I could explain the idea and remind you how to do it. Hope that helps =]